# Adaptive Step Duration for Precise Foot Placement: Achieving Robust Bipedal Locomotion on Terrains with Restricted Footholds

### v2

@@ V-A Robust Walking on Stepping Stones
The stepping stone profiles are listed in Table I. Each profile contains 32 stones with each relative position computed using $(L^{j},W^{j})$ as $\mathbf{p}_{stone,\mathrm{rel}}^{j}=[L^{j},w_{l/r}+W^{j}]^{T}$ ($j=\{1,...,32\}$), where $w_{l/r}$ is chosen as ±$\pm 0.28$ meters.
In the simulation, the robot is initialized in a standing posture with zero velocity and then walks 6 steps to approach the first stepping stone. In this speeding-up stage, the robot is only commanded to reach an appropriate $x$-velocity before walking onto the stepping stones with a right swing phase and thus may run into different initial conditions for the stepping stone test. Though such uncertainty helps highlight the robustness of our planner, we try to ensure consistent initial conditions in each comparison test between ALIP-MPC and ALIP-QP. As we will discuss below, ALIP-MPC outperforms ALIP-QP in all the presented simulation results.
@@ V-A1 Profile I
Table II lists the step position errors that are computed as the root-mean-square errors (RMSE) between the actual step positions and stone positions. For the test of $L^{j}=0.2$ and $0.4$ meters, ALIP-MPC slightly reduces the step position errors compared to ALIP-QP. This is because these tests all lead to periodic gaits and thus the multi-step preview provides similar information as the one-step preview. However, in the test of $L^{j}=0.55$ meters, the desired step length becomes too large and close to the mechanical limit. ALIP-QP fails to maintain a viable gait starting from the first step onto the stepping stone due to the harsh initial condition, while ALIP-MPC completes all the tests thanks to the strategy that considers the viability of multiple future steps to decide the next foot placement.
TABLE II: Step Position Errors in Profile I
$W^{j}$ $L^{j}$ $0.2$ $0.4$ $0.5$
$0.15$ $\mathbf{0.036}$ $0.040$ $\mathbf{0.020}$ $0.029$ $\mathbf{0.028}$ N/A
$-0.15$ $\mathbf{0.039}$ $0.042$ $\mathbf{0.019}$ $0.024$ $\mathbf{0.022}$ N/A
  • Note: Comparison of step position errors (in meters) in the test of Profile I. ALIP-MPC outperforms ALIP-QP in these tests, where ALIP-QP fails in the tests of large step lengths with $L^{j}=0.55$ (m).
@@ V-A2 Profile II
Each relative stone position is given by randomly generated $(L^{j},W^{j})$. ALIP-MPC completes this test with a step position error of $0.023$ meters, while ALIP-QP ends halfway with the robot falling. Fig. 4 compares the footprints, the CoM, and the DCM trajectories of two planners. ALIP-QP fails at the 14th step because the 13th foot placement yields an initial condition that cannot lead to viable states given the 14th stone position, highlighting the significance of the multi-step preview. Fig. 5 compares the step duration and initial DCM in $y$-direction of two planners, where ALIP-MPC can find foot placements to maintain viable states with smoother changes in the actual step duration.
Figure 4: Comparison of footprints in Profile II test. The black dots are the actual step positions, with the black frames illustrating the flat feet of the Digit, while the green rectangles mark the step position bounds on the gray stepping stones. (Left, ALIP-MPC) The robot walks stably through the random stones with a bounded DCM; (Right, ALIP-QP) The robot falls at the 14th step due to a non-viable foot placement of the 13th step (marked as red frames).
Figure 5: Comparison of step duration and initial DCM in the test of Profile II. The vertical gray lines mark the touch-down moments of walking steps. (Left, ALIP-MPC) The step duration is adjusted to bound the DCM with smoother changes; (Right, ALIP-QP) ALIP-QP fails to bound the DCM.
@@ V-A3 Profile III and IV
These two profiles emphasize on the abrupt changes in the $L^{j}$ and $W^{j}$, respectively. The step position errors of ALIP-MPC in these two tests are $0.027$ and $0.024$ meters, while ALIP-QP falls after a few steps. Fig. 6 shows the footprints of ALIP-MPC in both tests. Even under such challenging stepping stone profiles, especially with large variations in stone positions in the lateral direction, ALIP-MPC can find foot placements to maintain viable states.
Figure 6: Footprints of ALIP-MPC in the tests of Profile III (Upper) and Profile IV (Lower). The robot completes these tests without any step violating the step position constraint.
@@ V-B Perturbation Test on Stepping Stones
Figure 7: Robot states of ALIP-MPC in the perturbation test. Four perturbation forces are marked by red, yellow, green, and purple arrows, respectively. (Upper Left) The $x$- and $y$-velocity changes abruptly due to the perturbation but is quickly restored; (Upper Right) The step duration is adjusted in response to the perturbation forces to maintain viable states; (Lower Left and Right) The initial DCM in both $x$- and $y$-direction are bounded.
This test scenario contains 24 steps on the stepping stones profile $(L^{j},W^{j})=(0.4,0)$ meters and four force perturbations applied from all four directions to the pelvis of the Digit robot. Each perturbation force lasts from the relative time $t=0.1$ to $0.2$ seconds in the specified step with the forces and directions as follows: 1) 5th step, $F_{x}=150$ N; 2) 10th step, $F_{x}=-150$ N; 3) 15th step, $F_{y}=75$ N; 4) 20th step, $F_{y}=-75$ N. The robot states during the perturbation test are plotted in Fig. 7. Since we choose a stone profile where the robot walks straightforwardly, the perturbation resistance of both planners in the $x$-direction is similar. When the first perturbation force pushes the robot forward, the robot takes the next step much faster than usual to keep balance (with increased CoM velocity and decreased step duration). The second push in the backward direction causes the robot to slow down and thus increases the step duration. In both tests, our approach demonstrates excellent capability of handling adversarial disturbances. When the push directions are orthogonal (i.e., $y$ direction) to the walking directions, the feasibility of viable steps is significantly reduced. The baseline ALIP-QP approach fails under disturbance, whereas the ALIP-MPC remains stable. This is because ALIP-MPC considers a longer horizon of locomotion, and therefore, is capable of regulating the durations of multiple future steps to improve stability and robustness.
@@ VI Conclusion
This paper presents a discrete-time MPC formulation for multi-step preview foot placement planning with variable step duration. We demonstrate its performance of achieving more robust robotic walking on challenging terrains. Compared to other MPC formulations for stepping stones, the discrete dynamics of the initial DCM vastly simplify the optimization problem. However, due to the bilinear constraint in the optimization constraints, the average solving time of each iteration is around $30$ to $50$ ms. Future work will focus on improving the computational efficiency of the proposed approach and validating its effectiveness in real-world experiments. We will further develop more rigorous criteria for gait robustness on complex terrains.

### v3

@@ V-A Robust Walking on Stepping Stones
The stepping stone profiles are listed in Table I. Each profile contains 32 stones with each relative position computed using $(L_{j},W_{j})$ as $\mathbf{p}_{stone,\mathrm{rel},j}=[L_{j},w^{l/r}+W_{j}]^{T}$ ($j=\{1,...,32\}$), where $w^{l/r}$ is chosen as $\pm 0.28$ meters.
In the simulation, the robot is initialized in a standing posture with zero velocity and then walks 6 steps to approach the first stepping stone. In this speeding-up stage, the robot is only commanded to reach an appropriate $x$-velocity before walking onto the stepping stones and thus may run into different initial conditions for the stepping stone test. Though such uncertainty helps highlight the robustness of our planner, we try to ensure consistent initial conditions in each comparison test between DCM-MPC and DCM-QP. As we will discuss below, DCM-MPC outperforms DCM-QP in all the presented simulation results.
TABLE II: Step Position Errors in Profile I
$W_{j}$ $L_{j}$ $0.2$ $0.4$ $0.6$
$0.15$ $\mathbf{0.011}$ $0.018$ $\mathbf{0.021}$ 0.027 $\mathbf{0.025}$ N/A
0.15-0.15 $\mathbf{0.014}$ $0.034$ $\mathbf{0.028}$ N/A $\mathbf{0.030}$ N/A
  • Comparison of step position errors (in meters) in the test of Profile I (left and bold: DCM-MPC, right: DCM-QP). DCM-MPC outperforms DCM-QP in these tests, where “N/A” indicates that DCM-QP fails.
@@ V-A1 Profile I
Table IIlists the step position errors that are computed as the root-mean-square errors (RMSE) between the actual step positions and stone positions. DCM-MPC completes all the tests successfully with small step position errors, while DCM-QP is more likely to fail when $L_{j}$ takes large values such as $0.6$ meters (which is close to the mechanical limit of the robot). Note that the aforementioned initial conditions are not symmetric for the tests with $W_{j}=0.15$ and 0.15-0.15 meters, where the latter seems to be more challenging because the first stepping stone in the tests usually requires the right swing foot. DCM-QP fails to maintain a viable gait starting from the first step onto the stepping stone due to such harsh initial conditions, while DCM-MPC benefits from the strategy that considers the viability of multiple future steps to decide the next foot placement.
Figure 4: Comparison of footprints in Profile II test. The black triangles mark the actual step positions, with the black frames illustrating the flat feet of the Digit, while the green rectangles mark the step position bounds on the gray stepping stones. (Left, DCM-MPC) The robot walks stably through the random stones with the bounded DCM; (Right, DCM-QP) The robot falls after the 24th step (marked as red frames) due to a non-viable foot placement of the 23rd step.
Figure 5: Comparison of step duration and initial DCM in the test of Profile II. The vertical gray lines mark the touchdown moments of walking steps. (Left, DCM-MPC) The step duration is adjusted to bound the DCM with smoother changes; (Right, DCM-QP) DCM-QP fails with a boundary step duration that indicates the non-viable DCM.
Figure 6: Footprints of DCM-MPC in the tests of Profile III (Upper) and Profile IV (Lower). The robot completes these tests without any step violating the step position constraint.
Figure 7: Robot states of DCM-MPC in the perturbation test. Four perturbation forces are marked by red, yellow, green, and purple arrows, respectively. (Upper Left) The $x$- and $y$-velocity change abruptly due to the perturbations but are quickly restored; (Upper Right) The step duration is adjusted in response to the perturbations to maintain viable states; (Lower Left and Right) The initial DCM in both $x$- and $y$-directions are bounded.
@@ V-A2 Profile II
Each relative stone position is given by randomly generated $(L_{j},W_{j})$. DCM-MPC completes this test with a step position error of $0.029$ meters, while DCM-QP ends halfway with the robot falling. Fig. 4 compares the footprints, the CoM, and the DCM trajectories of two planners. DCM-QP fails to place its foot onto the desired 24th stone because the 23rd foot placement yields an initial DCM that cannot lead to viable states given the 24th stone position, highlighting the significance of the multi-step preview. Fig. 5 compares the step duration and initial DCM in $y$-direction of two planners, where DCM-MPC can find foot placements to maintain viable states with smoother changes in the actual step duration.
@@ V-A3 Profile III and IV
These two profiles emphasize the abrupt changes in the $L_{j}$ and $W_{j}$, respectively. The step position errors of DCM-MPC in these two tests are $0.025$ and $0.026$ meters, while DCM-QP falls after a few steps. Fig. 6 shows the footprints of DCM-MPC in both tests. Even under such challenging stepping stone profiles, especially with large variations in stone positions in the lateral direction, DCM-MPC can find foot placements to maintain viable states.
@@ V-B Perturbation Test on Stepping Stones
The test scenario with perturbations contains 24 steps on the stepping stones profile $(L_{j},W_{j})=(0.4,0)$ meters with the step width bound relaxed by $0.1$ meters and four force perturbations applied from all four directions to the pelvis of the Digit robot. Each perturbation force lasts from the relative time $t=0.1$ to $0.2$ seconds in the specified step with the forces and directions as follows: 1) 5th step, $F^{x}=175$ N; 2) 10th step, $F^{x}=-175$ N; 3) 15th step, $F^{y}=125$ N; 4) 20th step, $F^{y}=-125$ N. The robot states during the perturbation test are plotted in Fig. 7. Since we choose a stone profile where the robot walks straightforwardly, the perturbation resistance of both planners in the $x$-direction is similar. During the first two perturbations, our approach demonstrates excellent capability of handling adversarial disturbances, where the step duration of the current step is rapidly adjusted to keep balance. When perturbation directions are orthogonal to the walking direction (i.e., in the $y$-direction), the feasibility of viable steps is significantly reduced. The baseline DCM-QP approach fails under such disturbances, whereas the DCM-MPC remains stable by considering a longer horizon of locomotion, and hence capable of regulating the durations of multiple future steps to improve stability and robustness.
@@ VI Conclusion
This paper presents a discrete MPC formulation for multi-step preview foot placement planning based on the DCM and capturability analysis. We demonstrate its performance in achieving more robust bipedal robotic walking on challenging terrains with highly restricted footholds and severe perturbations, which benefits from variable step duration. Compared to other MPC formulations for stepping stones, the discrete dynamics of the initial DCM vastly simplify the optimization problem, thus ensuring real-time implementation. Future work will focus on validating its effectiveness in real-world experiments and further analyzing gait robustness on uneven terrains.